OpenStudy (anonymous):

SOMEONE PLEASE HELP! :( The midpoint of UV is (5,-11). The coordinates of one endpoint are U(3,5). Find the coordinates of endpoint V.

OpenStudy (anonymous):

\[\left( x_m,\,y_m \right)=\left( \frac{ x_1+x_2 }{ 2 },\,\frac{ y_1+y_2 }{2 } \right)\]

OpenStudy (anonymous):

Equate (3+x), (5+y) = (5, -11)

OpenStudy (anonymous):

Sorry Equate ( (3+x)/2), ((5+y)/2) = (5, -11)

OpenStudy (anonymous):

So do I just put the (5,-11) in for x and y?

OpenStudy (anonymous):

So (1.5+0.5x) = 5 (2.5+0.5y) = -11

OpenStudy (anonymous):

double and subtract... (10-3, -22 -5)

OpenStudy (anonymous):

therefore x = (5-1.5)/0.5 y = (-11-2.5)/0.5

OpenStudy (anonymous):

hope that was helpful

OpenStudy (anonymous):

Eh, I'm still a little confused how you got that answer .-. thank you though!

OpenStudy (anonymous):

sorry, correct but confusing. the midpoint is halfway between points, or the average of of two coordinates. so if you double the midpoint and subtract off one of the endpoints, you get the other. you have to do this for x and y separately.

OpenStudy (anonymous):

so double 5 and -11 and subtract 3 to the doubled number of 5 and 5 to the doubled number of -11?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

what do you get?

OpenStudy (anonymous):

7, -27 ?

OpenStudy (anonymous):

me too!

OpenStudy (anonymous):

alright will you help me with two more? I'm not very good at geometry and I need all the help I can get!